If two sources of sound at frequencies "f1" and "f2" are present, a listener may hear a beat frequency at f_beat = |f1 - f2|. Say your lab instructor turns on two such sources, and tells you f1 = 500 Hz, and that f2 is slightly higher frequency. You hear the combined sound level from the sources get louder and softer, and you measure that it takes 10 seconds for this cycle to repeat 20 times. What is the frequency (in Hz) of the second source, f2? (Enter just the number, without units.)
Answer #1

given beat frequeny

f2-f1 = 20/10

since f2 is greater than f1

we have

f2-f1 = 2

f2 = f1 + 2

f2 = 502 Hz